OpenStudy (anonymous):

Surface area and volume of a cubic section using spherical coordinates:

OpenStudy (anonymous):

Do you have a specific problem in mind? What is it they you need to understand? You haven't provided enough detail.

OpenStudy (anonymous):

The question is to find the surface area of the spherical section and the enclosed volume of: \[0\le R \le2, 0 \le \theta \le 90^o, 30^o \le \phi \le 90^o\] and then draw it the section

OpenStudy (anonymous):

Is the integral for the surface area: \[\int\limits_{0}^{90^o} \int\limits_{30^o}^{90^o} R^2\sin \theta d \theta \]

OpenStudy (anonymous):

except you are missing \[d \phi \]

OpenStudy (anonymous):

Ok thanks what about the volume integral how would that look?

OpenStudy (anonymous):

try this wiki page for the details http://en.wikipedia.org/wiki/Spherical_coordinate_system The volume element and surface element is also listed under "Integration and differentiation in spherical coordinates"

OpenStudy (anonymous):

Thats great thanks for the help

OpenStudy (anonymous):

I suggest you review this http://tutorial.math.lamar.edu/Classes/CalcIII/TISphericalCoords.aspx this is a great site for calcIII stuff. Suggest referencing this when on shaky ground

OpenStudy (anonymous):

I use it frequently, one can't remember everything especially when I don't use stuff much like spherical coordinates. Cheers!

OpenStudy (anonymous):

Very true thanks again